# MoTCoder: Elevating Large Language Models with Modular of Thought for Challenging Programming Tasks

### v2

@@ 4 Experiments
We showcase the effectiveness of our MoTCoder in addressing challenging code generation tasks, specifically focusing on the following prominent benchmarks:
@@
4.1 Examination of Modular Ability
In this section, we evaluate the modular capabilities of our model through both normal and modular inference, utilizing prompts as illustrated in Figure 5. For modular prompts, we instruct the model to utilize provided submodules to complete the final code. Conventional models lack this ability to leverage submodules, making the logical and modular aspects more challenging for the model.
Results are in Table 1. On one hand, our MoTCoder demonstrates a significant performance boost compared to models with the same parameters trained with traditional approaches, achieving a remarkable 12.9% improvement on APPs. On the other hand, when employing modular reasoning on models that have not undergone MoT training, there is a considerable decrease in performance. In contrast, our outcomes are minimally affected. This confirms the superior modular capabilities of our approach, which is crucial in large-scale code development environments, where developers often leverage and integrate code sub-modules to construct code frameworks, rather than crafting each line of code from scratch. MoTCoder enables the construction of complete code using only a standardized description of functions, unlocking the potential to generate extensive and complex code frameworks.
The results of MoTCoder also demonstrates that the absence of submodule inputs has a negligible impact on the results. This implies that our method autonomously learns to design submodules without requiring additional hints. In subsequent experiments, we use normal prompts during the inference process for MoTCoder and other methods to ensure a fair comparison.
@@ 4.2 Ablation Experiments
In this section, we perform ablation experiments to examine the effectiveness of our MoTCoder training framework.
4.2.1 Modular-of-Thought Instruction-Tuning
We propose two instruction-tuning strategies:
  1. 1.
    One-step strategy. The models directly generate the necessary modules and incorporate them into a complete solution.
  2. 2.
    Two-step strategy. Initially, the instructions outline the essential sub-modules, generating their function headers and docstrings that describe their intended purpose. Subsequently, the instruction guides the model to implement these outlined modules and eventually merge them into a cohesive final solution.
Note that both strategies generate substeps within a single process. As indicated in Table 2, the suggested two-step modular-of-thought instruction tuning strategy outperforms the other both for normal and modular inference. This aligns with the problem-solving approach employed by adept developers: When faced with a problem, skilled developers design solutions with a modular structure, organizing them into high-level logical sub-modules. They subsequently finalize and integrate these components to effectively produce their final solutions.
4.2.2 Training Dataset
In this section, we conduct ablation experiments on the dataset, as illustrated in Table 3. Initially, we finetuned our model on the modular dataset obtained after modular-of-though instruction evolution, resulting in a substantial improvement (11.0%) compared to the baseline of normal fine-tuning. Subsequently, we explore the incorporation of a normal dataset, conducting experiments on a mixture of normal and modular datasets, revealing further performance enhancement of 3.38%. This underscores the significance of data diversity: While we encourage the model to decompose tasks into submodules, we also allow the model to employ direct methods to solve problems. This provides a larger solution space, leading to significant improvements, especially for easy and medium-level problems (2.46% and 4.54%, respectively). The improvement is more modest for difficult problems (0.84%), aligning with the intuition that direct methods are better suited for simpler problems.
Table 6: CodeContests test and validation results by pass@k (%). The rounds of revision after the first round of direct generation is 5 for CodeChain (Le et al., 2023). SOTA models includes code-davinci, GPT3.5 and WizardCoder (Luo et al., 2023b), with code-revision method CodeChain (Le et al., 2023) and filtering method CoT(Chen et al., 2023).
Model Size Revision Validation Test Average
pass@1 pass@5 pass@1 pass@5 pass@1 pass@5
code-davinci-002 - - - - 1.00 - 1.00 -
code-davinci-002 + CodeT - - - - 3.20 - 3.20 -
WizardCoder 15B - 1.11 3.18 1.98 3.27 1.55 3.23
WizardCoder + CodeChain 15B 5 2.35 3.29 2.48 3.30 2.42 3.30
MoTCoder 15B - 2.39 7.69 6.18 12.73 4.29 10.21
GPT3.5 - - 6.81 16.23 5.82 11.16 6.32 13.70
@@ 4.3 Results
In this section, we present a comprehensive comparison of our approach with existing large coding model baselines on APPS  (Hendrycks et al., 2021a) in Section 4.3.1 and CodeContest (Li et al., 2022a) in Section 4.3.2. We employ normal inference prompts for all methods.
@@ 4.3.1 Results on APPS
We conducted a comparison of our approach with existing large language model baselines on APPs (Hendrycks et al., 2021a). All outcomes are computed using raw predictions without being filtered by the test cases provided in the prompt. Our analysis includes a comparison with open-sourced approaches such as CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023), and one-shot WizardCoder (Luo et al., 2023b). Additionally, we present results from closed-source models, including text-davinci, code-davinci, and GPT3.5.
As depicted in the results from the APPS shown in Table 4, we notice that the modular-of-thought training approach leads to an enhancement in code generation capabilities compared with previous instruction finetuning models. Our MoTCoder exhibits improved performance across all difficulty levels and demonstrates more substantial gains in the interview and competition-level problems, characterized by more intricate solutions. To provide specifics, the pass@1 performance of MoTCoder-15B surpasses WizardCoder-15b by 15.49% on interview-level problems and 10.28% on competition-level problems. Moreover, our pass@1 on competition-level questions surprisingly outperformed the closed-source model GPT3.5 by an impressive margin of 5.7%. This outcome serves as compelling evidence of the efficacy of our approach in addressing intricate problems.
We also conduct a comparative analysis of our approach against previous LLM baselines with code-revision methods as well. Our included baselines contain Codex (Chen et al., 2021c), CodeT5 (Wang et al., 2021), code-davinci, StarCoder (Li et al., 2023), and WizardCoder (Luo et al., 2023b) and code-revision methods contain Self-edit (Zhang et al., 2023), CodeRL (Wang et al., 2021; Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023). The results presented in Table 5 illustrate that MoTCoder exhibits notable performance improvements.
Specifically, MoTCoder achieves a 20.80% pass@1, surpassing the current SOTA, CodeChain + WizardCoder, by 10.3%. CodeChain (Le et al., 2023) is an iterative framework for inference that elicits modularized code generation through a chain of self-revisions. In contrast to CodeChain, our approach introduces a guided modular-of-thought framework during training, making it more intrinsic. This leads to improvements in performance. Notably, this achievement is obtained without the need for any additional revision iterations, thereby avoiding any extra associated inference costs.
@@ 4.3.2 Results on CodeContests
We conduct a comprehensive evaluation of our approach on CodeContests, benchmarking it against current state-of-the-art models, including code-davinci, GPT3.5, and WizardCoder (Luo et al., 2023b). The competitive models utilize the code-revision method CodeChain (Le et al., 2023) and the filtering method CodeT (Chen et al., 2023). The results, as depicted in Table 6, reveal notable performance enhancements achieved by MoTCoder. Specifically, MoTCoder attains an impressive 12.73% pass@5 on the test set, surpassing WizardCoder with CodeChain by 9.43%. Remarkably, MoTCoder even outperforms GPT3.5 by 1.57%. Importantly, these achievements are realized without the need for additional revision iterations or filtering by test samples, thereby avoiding any additional associated inference costs or the introduction of extra information.
@@ 5 Limitations
Through experimentation, it has been observed that although our MoTCoder elevates the model’s modular and logical capabilities, leading to a commendable performance in addressing competitive-level programming challenges, it concurrently sacrifices some of the model’s conversational abilities. This finding, along with corroborative evidence from prior research (Roziere et al., 2023), underscores the persistent and crucial nature of addressing the forgetting problem. We remain committed to ongoing exploration of this issue in our future work.
@@ 6 Conclusion
We propose a novel framework, MoTCoder, for tuning instructions based on a modular of thought. MoTCoder directs LLMs to decompose their solutions into modular segments, culminating in the synthesis of a comprehensive final solution. Our experiments demonstrate that modular-of-thought strategy significantly improves LLM performance, achieving SOTA passing rate on challenging code tasks and showcasing superior modular capabilities. It is crucial in large-scale code development environments, since developers commonly leverage code sub-modules to construct code frameworks over crafting each line of code from scratch. MoTCoder empowers the creation of entire code structures using only standardized descriptions of functions. Adopting this methodology, we unlock the potential to generate extensive and intricate code frameworks.

### v3

@@ 4 Experiments
We demonstrate the efficacy of MoTCoder in tackling intricate code-generation tasks, rather than those that can be solved with just a few lines, as exemplified by HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021). Specifically, we focus on the following prominent benchmarks.
@@
Table 1: Training Datasets. We compare MoTCoder training on normal and modular-of-thought dataset. APPs test results by pass@1 (%) are shown.
Model Introductory Interview Competition All
Normal 50.20 26.65 11.52 28.33
Module-of-Thought 50.01 29.81 14.36 30.76
Table 2: APPS test results by pass@k (%). Competitive approaches include CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023b), WizardCoder (Luo et al., 2023c), CodeLlama series (Roziere et al., 2023), text-davinci, GPT-4 (OpenAI, 2023b). We also include code-revision methods containing Self-edit (Zhang et al., 2023), CodeRL (Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023).
Model Size Introductory Interview Competition All
CodeT5 770M 6.60 1.03 0.30 2.00
CodeRL+CodeT5 770M 7.08 1.86 0.75 2.69
text-davinci-002 - - - - 7.48
Self-edit+text-davinci-002 - - - - 7.94
GPT-2 0.1B 5.64 6.93 4.37 6.16
1.5B 7.40 9.11 5.05 7.96
GPT-Neo 2.7B 14.68 9.85 6.54 10.15
GPT-3 175B 0.57 0.65 0.21 0.55
StarCoder 15B 7.25 6.89 4.08 6.40
WizardCoder 15B 26.04 4.21 0.81 7.90
CodeChain+WizardCoder 15B 26.29 7.49 3.75 10.50
Octocoder 16B 16.50 7.92 4.61 8.97
Codellama 7B 14.15 6.63 4.00 7.61
13B 23.94 13.50 9.80 14.85
34B 32.01 18.61 10.19 19.61
Codellama-Python 7B 18.83 8.62 4.47 9.83
13B 26.40 13.44 6.86 14.72
34B 26.45 16.61 8.77 17.01
Codellama-Instruct 7B 14.20 6.63 4.43 7.70
13B 22.41 14.34 6.62 15.21
34B 28.64 16.80 10.51 17.91
Deepseek-Coder-Base 6.7B 40.23 22.12 13.04 23.92
Deepseek-Coder-Instruct 6.7B 44.65 23.86 12.89 25.83
MoTCoder 6.7B 50.01 29.81 14.36 30.76
GPT-4 - 34.97 13.75 14.52 18.15
@@ 4.1 Ablation Experiments
In this section, we conduct ablation experiments on the dataset, as illustrated in Table 1. We evaluated the performance of instruction tuning on our proposed Module-of-Thought (MoT) dataset versus original dataset, which is a small part of training set of CodeContests and APPS. We utilize the DeepSeek-Coder-6.7B model as the base model. The results, as delineated in the table below, highlight a pivotal finding: while the MoT tuning strategy maintains comparable performance on introductory tasks, it significantly enhances outcomes in competitive programming scenarios, elevating the pass@1 rate from 11.52% to 14.36%. The results align with our intuition that breaking down code into sub-modules is more suitable for complex tasks, whereas simpler tasks are better suited for direct generation. The overall enhancement underlines the effectiveness of our MoT method in elevating model proficiency in competitive programming domains.
Table 3: CodeContests test results by pass@k (%). The rounds of revision after the first round of direct generation is 5 for CodeChain Le et al. (2023). SOTA models include StarCoder Li et al. (2023b), WizardCoder Luo et al. (2023c), CodeLlama series models Roziere et al. (2023), code-davinci, GPT4 OpenAI (2023b), with code-revision method CodeChain Le et al. (2023) and filtering method CoTChen et al. (2023).
Model Size pass@1 pass@5
code-davinci-002 - 1.00 -
code-davinci-002 + CodeT - 3.20 -
WizardCoder 15B 1.98 3.27
WizardCoder + CodeChain 15B 2.48 3.30
Octocoder 16B 4.95 13.03
Codellama 7B 0.30 1.11
13B 2.12 6.26
34B 5.35 12.02
Codellama-Python 7B 4.75 10.30
13B 4.75 12.32
34B 5.86 14.85
Codellama-Instruct 7B 2.12 6.26
13B 5.96 12.02
34B 6.46 14.24
Deepseek-Coder-Base 6.7B 6.46 15.25
Deepseek-Coder-Instruct 6.7B 6.87 8.18
MoTCoder 6.7B 9.29 16.97
GPT-4 - 16.36 -
@@ 4.2 Results
In this section, we present a comprehensive comparison with existing large coding model baselines on APPS  (Hendrycks et al., 2021a) and CodeContest (Li et al., 2022a).
@@ 4.2.1 Results on APPS
We conducted a comparison of our approach with existing large language model baselines on APPs (Hendrycks et al., 2021a). All outcomes are computed using raw predictions without being filtered by the test cases provided in the prompt. Our analysis includes a comparison with open-sourced approaches such as CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder Li et al. (2023b), WizardCoder (Luo et al., 2023c), and CodeLlama series (Roziere et al., 2023). Additionally, we present results from SOTA closed-source models GPT-4 (OpenAI, 2023b).
As depicted in the results from the APPS shown in Table 2, we notice that the module-of-thought training approach leads to an enhancement in code generation capabilities compared with previous instruction finetuning models. Our MoTCoder exhibits improved performance across all difficulty levels and demonstrates more substantial gains in the interview and competition-level problems, characterized by more intricate solutions. To provide specifics, the pass@1 performance of MoTCoder-6.7B surprisingly outperformed the closed-source model GPT-4 by an impressive margin of 12.61%. This outcome serves as compelling evidence of the efficacy of our approach in addressing competitive programming problems.
We also conduct a comparative analysis of our approach against previous LLM baselines with code-revision methods as well. Our included baselines contain Codex (Chen et al., 2021c), CodeT5 (Wang et al., 2021), code-davinci, StarCoder (Li et al., 2023b), and WizardCoder (Luo et al., 2023c) and code-revision methods contain Self-edit (Zhang et al., 2023), CodeRL (Wang et al., 2021; Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023). The results presented in Table 2 illustrate that MoTCoder exhibits notable performance improvements.
Specifically, MoTCoder achieves a 30.76% pass@1, surpassing CodeChain + WizardCoder, by 20.26%. CodeChain (Le et al., 2023) is an iterative framework for inference that elicits modularized code generation through a chain of self-revisions. In contrast to CodeChain, our approach introduces a guided module-of-thought framework during training, making it more intrinsic. This leads to improvements in performance. Notably, this achievement is obtained without the need for any additional revision iterations, thereby avoiding any extra inference costs.
@@ 4.2.2 Results on CodeContests
We conduct an evaluation of our approach on CodeContests (Li et al., 2022a), benchmarking it against current state-of-the-art models, including code-davinci, GPT3.5, and WizardCoder (Luo et al., 2023c). The competitive models utilize the code-revision method CodeChain (Le et al., 2023) and the filtering method CodeT (Chen et al., 2023). The results, as depicted in Table 3, reveal notable performance enhancements achieved by MoTCoder. Specifically, MoTCoder attains an impressive 12.73% pass@5 on the test set, surpassing WizardCoder with CodeChain by 9.43%. These achievements are realized without the need for additional revision iterations or filtering by test samples, thereby avoiding any additional associated inference costs or the introduction of extra information.
@@ 5 Limitations
Through experimentation, it has been observed that although our MoTCoder elevates the model’s modular and logical capabilities, leading to a commendable performance in addressing competitive-level programming challenges, it concurrently sacrifices some of the model’s conversational abilities. This finding, along with corroborative evidence from prior research Roziere et al. (2023), underscores the persistent and crucial nature of addressing the forgetting problem. We remain committed to ongoing exploration of this issue in our future work.
@@ 6 Conclusion
This study highlights the limitations of Large Language Models (LLMs) in solving complex programming tasks due to their tendency to generate monolithic code blocks. In response, we developed Module-of-Thought Coder (MoTCoder), a novel framework that encourages the breakdown of tasks into manageable sub-tasks and sub-modules. Our results demonstrate that MoTCoder’s approach significantly enhances the modularity and accuracy of solutions, as evidenced by considerable improvements in pass@1 rates on both APPS and CodeContests benchmarks. We believe the introduction of MoT instruction tuning as a method to cultivate and leverage sub-modules paves the path for a promising direction for future research.
Appendix A Broader impacts
Positive societal impacts.
We propose a method for challenging code programming, which will be helpful for users to solve code problems.
negative societal impacts.
This approach may be abused in programming contests that are supposed to be done by humans.
Appendix B Examples
For given input problems, we provided submodules and final code solution examples for problems from APPs test dataset.
B.0.1 APPs Example 1
Listing 1: Problem 1 from APPs test dataset
An accordion is a string (yes, in the real world accordions are musical instruments, but lets forget about it for a while) which can be represented as a concatenation of: an opening bracket (ASCII code \$091\$), a colon (ASCII code \$058\$), some (possibly zero) vertical line characters (ASCII code \$124\$), another colon, and a closing bracket (ASCII code \$093\$). The length of the accordion is the number of characters in it.
For example, [::], [:||:] and [:|||:] are accordions having length \$4\$, \$6\$ and \$7\$. (:|:), {:||:}, [:], ]:||:[ are not accordions.
You are given a string \$s\$. You want to transform it into an accordion by removing some (possibly zero) characters from it. Note that you may not insert new characters or reorder existing ones. Is it possible to obtain an accordion by removing characters from \$s\$, and if so, what is the maximum possible length of the result?
—–Input—–
The only line contains one string \$s\$ (\$1 $\le$ |s| $\le$ 500000\$). It consists of lowercase Latin letters and characters [, ], : and |.
—–Output—–
If it is not possible to obtain an accordion by removing some characters from \$s\$, print \$-1\$. Otherwise print maximum possible length of the resulting accordion.
—–Examples—–
Input
|[a:b:|]
Output
4
Input
Output
-1
Listing 2: Submodules generated for Problem 1 from APPs test dataset
def find_colon_indices(s):
Description: This function finds the indices of the first and last occurrence of ":" in the given string.
Input:
s (str): A string containing colons.
Output:
ind (int): The index of the first occurrence of ":".
bind (int): The index of the last occurrence of ":".
def count_pipes_between_colons(s, ind, bind):
Description: This function counts the number of "|" characters between the first and last occurrence of ":" in the given string.
Input:
s (str): A string containing colons and "|" characters.
ind (int): The index of the first occurrence of ":".
bind (int): The index of the last occurrence of ":".
Output:
ans (int): The count of "|" characters between the first and last occurrence of ":".
Listing 3: Code solution generated for Problem 1 from APPs test dataset
def find_colon_indices(s):
ind = s.find(":")
bind = s.rfind(":")
return ind, bind
def count_pipes_between_colons(s, ind, bind):
ans = 0
if ind!= 0:
ans += s.count("|", 0, ind)
if bind!= len(s) - 1:
ans += s.count("|", bind + 1, len(s))
return ans
s = input()
ind, bind = find_colon_indices(s)
ans = count_pipes_between_colons(s, ind, bind)
if ind == -1 or bind == -1 or ind > bind:
print(-1)
else:
print(bind - ind + 1 + ans)
B.0.2 APPs Example 2
Listing 4: Problem 2 from APPs test dataset
Anton has the integer x. He is interested what positive integer, which doesnt exceed x, has the maximum sum of digits.
Your task is to help Anton and to find the integer that interests him. If there are several such integers, determine the biggest of them.
—–Input—–
The first line contains the positive integer x \$(1 $\leq$ x $\leq$ 10^18)\$ - the integer which Anton has.
—–Output—–
Print the positive integer which doesnt exceed x and has the maximum sum of digits. If there are several such integers, print the biggest of them. Printed integer must not contain leading zeros.
—–Examples—–
Input
100
Output
99
Input
48
Output
48
Input
521
Output
499
Listing 5: Submodules generated for Problem 2 from APPs test dataset
def find_best_number(num):
Description: This function finds the best number by decrementing digits from the given number.
Input:
num (list): A list of digits representing the number.
Output:
best (list): A list of digits representing the best number.
def convert_to_string(num):
Description: This function converts a list of digits into a string representation.
Input:
num (list): A list of digits.
Output:
s (str): The string representation of the digits.
Listing 6: Code solution generated for Problem 2 from APPs test dataset
def find_best_number(num):
best = num[:]
for i in range(8, -1, -1):
if num[i] > 0:
num[i] -= 1
break
return num
def convert_to_string(num):
s = .join(map(str, num))
return s
x = int(input())
num = [int(i) for i in str(x)]
best = find_best_number(num)
s = convert_to_string(best)
print(s)
print(s)